OpenStudy (anonymous):

What of the following statements have the same result? f(1) when f(x) = 5x + 1 f−1(3) when f(x) = 2x+3 3y − 7 = y + 5

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(x) = 5x + 1 f(1) = 5(1) + 1 ... replace every x with 1 f(1) = ???

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

close but no

OpenStudy (anonymous):

6 lol

jimthompson5910 (jim_thompson5910):

"5(1) + 1" means "5 times 1 plus 1"

jimthompson5910 (jim_thompson5910):

yeah f(1) = 6

jimthompson5910 (jim_thompson5910):

"f−1(3) when f(x) = 2x+3" is a fancy way of saying "plug in f(x) = 3 and solve for x" f(x) = 2x+3 3 = 2x+3 3-3 = 2x+3-3 ... Subtract 3 from both sides. 0 = 2x x = ???

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

yep x = 0 leads to f(x) = 3

jimthompson5910 (jim_thompson5910):

solve 3y − 7 = y + 5 for y. Tell me what you get

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

y = 6 is correct

jimthompson5910 (jim_thompson5910):

from part a) we got f(1) = 6

OpenStudy (anonymous):

ok so 1 and 3 :)

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

I have one last one and I'm done for tonight!

jimthompson5910 (jim_thompson5910):

show me what you have so far

OpenStudy (anonymous):

Ok. so I guess a positive end behavior for f(x)+2

OpenStudy (anonymous):

yes?

jimthompson5910 (jim_thompson5910):

how do you know it's positive?

OpenStudy (anonymous):

i just assumbed bc the leading co. is positive

jimthompson5910 (jim_thompson5910):

we don't know anything about the original function f(x). So we can't say if it has positive or negative end behavior

jimthompson5910 (jim_thompson5910):

what we can say is that the end behavior won't change if we add 2 to f(x)

OpenStudy (anonymous):

true. I agree

jimthompson5910 (jim_thompson5910):

f(x)+2 just shifts f(x) up 2 units

jimthompson5910 (jim_thompson5910):

on the other hand (-1/2)*f(x) flips f(x) over the x axis and compresses it vertically by a factor of 2

jimthompson5910 (jim_thompson5910):

so (-1/2)*f(x) will have its end behavior flipped

OpenStudy (anonymous):

I agree.

jimthompson5910 (jim_thompson5910):

what can you say about the y-intercept?

OpenStudy (anonymous):

It would be (0,2)

jimthompson5910 (jim_thompson5910):

but do we know what the y-intercept of f(x) is?

OpenStudy (anonymous):

and increasing?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

we don't know what the y-intercept of f(x) is

jimthompson5910 (jim_thompson5910):

but whatever it is, it is shifted up 2 units for f(x)+2

OpenStudy (anonymous):

yes. And what about the -1/2 one?

OpenStudy (anonymous):

so a wide parabola?

jimthompson5910 (jim_thompson5910):

it gets wider when you compress it vertically, yes

OpenStudy (anonymous):

But I don't understand how it is increasing and the regions where it does part>>

OpenStudy (anonymous):

any ideas?

jimthompson5910 (jim_thompson5910):

f(x) isn't given, so we cannot find the increasing/decreasing intervals. Whatever they are, they don't change when going to f(x)+2. Everything shifts up which is why the intervals don't change

jimthompson5910 (jim_thompson5910):

with (-1/2)*f(x), the intervals swap. Whatever was decreasing is now increasing and vice versa.

OpenStudy (anonymous):

Oh:O !! That makes sooo much more sense! Thanks:)

OpenStudy (anonymous):

is that it? We are done with the question

jimthompson5910 (jim_thompson5910):

yeah I think so